# Descending finite projective modules from a Novikov ring

### v1

@@ 3. Novikov isocrystals
A Novikov isocrystal over a ring $A$ is a pair $(M,F_{M})$, where $M$ is a finite projective $A(\!(t^{\mathbb{R}})\!)$-module and $F_{M}\colon M\to M$ is a $F$-semilinear bijection. A morphism $f\colon(M,F_{M})\to(N,F_{N})$ between two Novikov isocrystals is a $A(\!(t^{\mathbb{R}})\!)$-linear map $f\colon M\to N$ satisfying $f\circ F_{M}=F_{N}\circ f$. We denote by $\mathsf{NovIsoc}(A)$ the category of Novikov isocrystals over $A$.
@@
Let $M_{0}$ and $N_{0}$ be two finite projective $A$-modules, and let $f\colon M_{0}\otimes_{A}A(\!(t^{\mathbb{R}})\!)\to N_{0}\otimes_{A}A(\!(t^{\mathbb{R}})\!)$ be a $A(\!(t^{\mathbb{R}})\!)$-linear map satisfying $f\circ(\operatorname{id}_{M_{0}}\otimes F)=(\operatorname{id}_{N_{0}}\otimes F)\circ f$. Regarding $f$ as an element of
@@
we see that $f$ must be fixed under the automorphism $\operatorname{id}_{\operatorname{Hom}_{A}(M_{0},N_{0})}\otimes F$, and hence $\operatorname{supp}(f)=\lambda\operatorname{supp}(f)$. On the other hand, the intersection $[-1,1]\cap\operatorname{supp}(f)$ is finite, and therefore $\operatorname{supp}(f)\subseteq\{0\}$. This shows that $f\in\operatorname{Hom}_{A}(M_{0},N_{0})$. ∎
@@ 3.1. Some results on isocrystals
We imitate the argument in the proof of the Dieudonné–Manin classification [Man63]. Because $k(\!(t^{\mathbb{R}})\!)$ is a field, it suffices to prove that for every nonzero Novikov isocrystal $(M,F_{M})$, there exists a nonzero element $m\in M$ and $c\in k^{\times}$ satisfying $F_{M}(m)=cm$. Fix an arbitrary nonzero $m\in M$ and find $n\geq 1$ such that $m,F_{M}(m),\dotsc,F_{M}^{n-1}(m)$ are linearly independent while
@@
for $a_{1},\dotsc,a_{n}\in k(\!(t^{\mathbb{R}})\!)$. Replacing $m$ with $t^{r}m$ for a real number $r\in\mathbb{R}$ has the effect of replacing $a_{i}$ with $t^{r(\lambda^{n}-\lambda^{i})}$. Thus by choosing $r$ appropriately, we may further assume that $a_{1},\dotsc,a_{n}\in k[[t^{\mathbb{R}}]]$ with $a_{i}\in k[[t^{\mathbb{R}}]]^{\times}$ for some $10\leq i\leq n-$ 1.
@@
We claim that this equation has a nonzero solution in $k[[t^{\mathbb{R}}]]$. Let $\bar{a}_{i}\in k$ be the constant terms of $a_{i}\in k[[t^{\mathbb{R}}]]$. Since we are assuming that $\bar{a}_{i}\neq 0$ for some $11\leq i\leq n-$ 1 and that $k$ is algebraically closed, there exists a $c\in k^{\times}$ for which
@@
$g(b)=c^{-n}F^{n}(b)F^{n-1}(a_{0})+c^{-n+1}F^{n-1}(b)F^{n-2}(a_{1})+\dotsb+c^{-1}F(b)$
so that we need to find $b\in k[[t^{\mathbb{R}}]]$ with $g(b)=b$. Set $b^{(0)}=1$ and $b^{(j+1)}=g(b^{(j)})$. By construction of $c$, the constant term of $b^{(1)}=g(b^{(0)})$ is $1$, and hence there exists a positive real number $\epsilon>0$ for which $b^{(1)}-b^{(0)}\in t^{\epsilon}k[[t^{\mathbb{R}}]]$. On the other hand, we observe that $x-y\in t^{d}k[[t^{\mathbb{R}}]]$ implies $g(x)-g(y)\in t^{\lambda d}k[[t^{\mathbb{R}}]]$. This shows that $b^{(j+1)}-b^{(j)}\in t^{\lambda^{i}\epsilon}k[[t^{\mathbb{R}}]]$ and therefore
@@
to regard $M$ as a $A(\!(t^{\mathbb{R}})\!)$-submodule of $N_{0}(\!(t^{\mathbb{R}})\!)$ that is stable under the action of $\operatorname{id}_{N_{0}}\otimes F$. Note that $A(\!(t^{\mathbb{R}})\!)\hookrightarrow B(\!(t^{\mathbb{R}})\!)$ is a closed subring with the subspace topology. By writing $M$ as a direct summand of a finite free module, we see that $M$ is also closed in $N_{0}(\!(t^{\mathbb{R}})\!)$, and moreover the canonical topology on $M$ agrees with the subspace topology from $N_{0}(\!(t^{\mathbb{R}})\!)$.
@@
This is a closed $A(\!(t^{\mathbb{R}})\!)$-submodule of $(A^{\oplus n})(\!(t^{\mathbb{R}})\!)$ that is stable under $F^{\oplus n}$. Lemma 3.6 implies that there exists a $A$-submodule $K_{0}\subseteq A^{\oplus n}$ for which $K=K_{0}(\!(t^{\mathbb{R}})\!)$. On the other hand, $K$ has the subspace topology from $A(\!(t^{\mathbb{R}})\!)^{\oplus n}$, and therefore the $t$-adic topology on $K_{0}(\!(t^{\mathbb{R}})\!)$ agrees with the canonical topology. Hence Lemma 2.10 shows that $K_{0}$ is a finite projective $A$-module. At this point, we have
@@ 4.1. Isocrystal from a descent datum
[…]
@@
[…]
Both rows are equalizers, and all horizontal maps are $A(\!(t^{\mathbb{R}})\!)$-linear while the vertical maps are $F$-semilinear, if we use the ring homomorphisms $\pi_{2}^{\ast}\colon A(\!(t^{\mathbb{R}})\!)\to A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$ and $\rho_{3}^{\ast}\colon A(\!(t^{\mathbb{R}})\!)\to A(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!)$ to give every module a $A(\!(t^{\mathbb{R}})\!)$-module structure. Therefore we obtain an $F$-semilinear bijection $F_{M}\colon M\to M$, and this defines the functor
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 4.2. The case of a reduced ring
with the proper that each associated graded $M_{i}/M_{i-1}$ is isomorphic to $(k(\!(t^{\mathbb{R}})\!),c_{i}F)$ for some $c_{i}\in k^{\times}$.
@@
has an $i$-dimensional kernel. Since the kernel when restricted to $\pi_{2}^{\ast}M_{i-1}$ is $(i-1)$-dimensional, there exist $x_{1},\dotsc,x_{i}\in k(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$ with $x_{i}\neq 0$ satisfying
@@
[…]
@@
For every element $\alpha=(\alpha_{i})\in M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$, define the support of $\alpha$ as $\operatorname{supp}(\alpha)=\bigcup_{i\in I}\operatorname{supp}(\alpha_{i})\subseteq\mathbb{R}$, where the support of $\alpha_{i}\in M_{0,i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$ is defined as before. For $\beta=(\beta_{i})\in M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}})\!)$ we similarly define $\operatorname{supp}(\beta)=\bigcup_{i}\operatorname{supp}(\beta_{i})\subseteq\mathbb{R}$.
@@
inside $\mathbb{R}^{2}$, where by $+$ we mean the Minkowski sum. By our assumptions, the right hand side is always contained in $\{(x,y):2x+y\geq C+D_{i}\}$. On the other hand, for some $1\leq j\leq n$, the left hand side contains $(D_{i},0)$. This shows that $2D_{i}\geq C+D_{i}$, and hence $D_{i}\geq C$. This shows that $\operatorname{supp}(m_{j,i})\subseteq[C,\infty)$ for all $i,j$, and taking the union over all $i$, we obtain $\operatorname{supp}(m_{j})\subseteq[C,\infty)$ for all $1\leq j\leq n$. For an arbitrary $m\in M$, we can now write $m=\sum_{j=1}^{n}a_{j}m_{j}$ for $a_{j}\in A(\!(t^{\mathbb{R}})\!)$, and since each $\operatorname{supp}(a_{j})$ is bounded below, we conclude that $\operatorname{supp}(m)$ is bounded below as well.
@@ 4.3. The case of a general ring
Since the Čech $H^{2}$ of quasi-coherent sheaves with respect to $\{U_{i}\}$ vanish on $\operatorname{Spec}A/I$ as each $U_{i}$ is affine, this is a coboundary, and hence we may modify the isomorphisms $M_{i}|_{U_{i}\cap U_{j}}\cong M_{j}|_{U_{i}\cap U_{j}}$ so that they do satisfy the cocycle condition. ∎
@@
$\operatorname{Hom}_{A(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!)}(M_{0}(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!),(IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!))$ $\cong\operatorname{Hom}_{A}(M_{0},IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!).$
Using this isomorphism, we identify $\psi_{+}$ as an element of $\operatorname{Hom}_{A}(M_{0},IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$, and similarly the additive cocycle condition as an equation valued in the group $\operatorname{Hom}_{A}(M_{0},IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!)$. By considering the coefficients in the cocycle condition, we readily see that $\psi_{+}$ is supported on ({0})×{0}{0}×($(\mathbb{R}\setminus\{0\})\times\{0\}\cup\{0\}\times(\mathbb{R}\setminus\{0\})$, and moreover the coefficient of $t^{d}$ agrees with the coefficient of $u^{d}$ for each $d\in\mathbb{R}\setminus\{0\}$. This allows us to write
@@
Regarding $\xi_{+}$ as a $A(\!(t^{\mathbb{R}})\!)$-linear map $M_{0}(\!(t^{\mathbb{R}})\!)\to(IM_{0})(\!(t^{\mathbb{R}})\!)$, we now define $\xi=\operatorname{id}+\xi_{+}$. It follows that

### v2

@@ 3. Novikov isocrystals
A Novikov isocrystal over a ring $A$ is a pair $(M,F_{M})$, where $M$ is a finite projective $A(\!(t^{\mathbb{R}})\!)$-module and $F_{M}\colon M\to M$ is a $F$-semilinear bijection. A morphism $f\colon(M,F_{M})\to(N,F_{N})$ between two Novikov isocrystals is an $A(\!(t^{\mathbb{R}})\!)$-linear map $f\colon M\to N$ satisfying $f\circ F_{M}=F_{N}\circ f$. We denote by $\mathsf{NovIsoc}(A)$ the category of Novikov isocrystals over $A$.
@@
Let $M_{0}$ and $N_{0}$ be two finite projective $A$-modules, and let $f\colon M_{0}\otimes_{A}A(\!(t^{\mathbb{R}})\!)\to N_{0}\otimes_{A}A(\!(t^{\mathbb{R}})\!)$ be an $A(\!(t^{\mathbb{R}})\!)$-linear map satisfying $f\circ(\operatorname{id}_{M_{0}}\otimes F)=(\operatorname{id}_{N_{0}}\otimes F)\circ f$. Regarding $f$ as an element of
@@
we see that $f$ must be fixed under the automorphism $\operatorname{id}_{\operatorname{Hom}_{A}(M_{0},N_{0})}\otimes F$, and hence $\operatorname{supp}(f)=\lambda\operatorname{supp}(f)$. On the other hand, the intersection $1,1]\cap\operatorname{supp}(f)$ is finite, and therefore $\operatorname{supp}(f)\subseteq\{0\}$. This shows that $f\in\operatorname{Hom}_{A}(M_{0},N_{0})$. ∎
@@ 3.1. Some results on isocrystals
We imitate the argument in the proof of the Dieudonné–Manin classification [Man63]. Because $k(\!(t^{\mathbb{R}})\!)$ is a field, it suffices to prove that for every nonzero Novikov isocrystal $(M,F_{M})$, there exists a nonzero element $m\in M$ and $c\in k^{\times}$ satisfying $F_{M}(m)=cm$. Fix an arbitrary nonzero $m\in M$ and find $n\geq 1$ such that $m,F_{M}(m),\dotsc,F_{M}^{n-1}(m)$ are linearly independent over $k(\!(t^{\mathbb{R}})\!)$ while
@@
for $a_{0},\dotsc,a_{n-1}\in k(\!(t^{\mathbb{R}})\!)$. Replacing $m$ with $t^{r}m$ for a real number $r\in\mathbb{R}$ has the effect of replacing $a_{i}$ with $t^{r(\lambda^{n}-\lambda^{i})}$. Thus by choosing $r$ appropriately, we may further assume that $a_{1},\dotsc,a_{n}\in k[[t^{\mathbb{R}}]]$ with $a_{i}\in k[[t^{\mathbb{R}}]]^{\times}$ for some $0\leq i\leq n-1$.
@@
We claim that this equation has a nonzero solution in $k[[t^{\mathbb{R}}]]$. For each $0\leq i\leq n-1$, let $\bar{a}_{i}\in k$ be the constant term of $a_{i}\in k[[t^{\mathbb{R}}]]$. Since we are assuming that $\bar{a}_{i}\neq 0$ for some $1\leq i\leq n-1$ and that $k$ is algebraically closed, there exists a $c\in k^{\times}$ for which
@@
$g(b)=c^{-n}F^{n}(b)F^{n-1}(a_{0})+c^{-n+1}F^{n-1}(b)F^{n-2}(a_{1})+\dotsb+c^{-1}F(b)a_{n-1}$
so that we need to find $b\in k[[t^{\mathbb{R}}]]$ with $g(b)=b$. Set $b^{(0)}=1$ and $b^{(j+1)}=g(b^{(j)})$. By construction of $c$, the constant term of $b^{(1)}=g(b^{(0)})$ is $1$, and hence there exists a positive real number $\epsilon>0$ for which $b^{(1)}-b^{(0)}\in t^{\epsilon}k[[t^{\mathbb{R}}]]$. On the other hand, we observe that $x-y\in t^{d}k[[t^{\mathbb{R}}]]$ implies $g(x)-g(y)\in t^{\lambda d}k[[t^{\mathbb{R}}]]$. This shows that $b^{(j+1)}-b^{(j)}\in t^{\lambda^{j}\epsilon}k[[t^{\mathbb{R}}]]$ and therefore
@@
to regard $M$ as an $A(\!(t^{\mathbb{R}})\!)$-submodule of $N_{0}(\!(t^{\mathbb{R}})\!)$ that is stable under the action of $\operatorname{id}_{N_{0}}\otimes F$. Note that $A(\!(t^{\mathbb{R}})\!)\hookrightarrow B(\!(t^{\mathbb{R}})\!)$ is a closed subring with the subspace topology. By writing $M$ as a direct summand of a finite free module, we see that $M$ is also closed in $N_{0}(\!(t^{\mathbb{R}})\!)$, and moreover the canonical topology on $M$ agrees with the subspace topology from $N_{0}(\!(t^{\mathbb{R}})\!)$.
@@
This is a closed $A(\!(t^{\mathbb{R}})\!)$-submodule of $(A^{\oplus n})(\!(t^{\mathbb{R}})\!)$ that is stable under $F^{\oplus n}$. Lemma 3.6 implies that there exists an $A$-submodule $K_{0}\subseteq A^{\oplus n}$ for which $K=K_{0}(\!(t^{\mathbb{R}})\!)$. On the other hand, $K$ has the subspace topology from $A(\!(t^{\mathbb{R}})\!)^{\oplus n}$, and therefore the $t$-adic topology on $K_{0}(\!(t^{\mathbb{R}})\!)$ agrees with the canonical topology. Hence Lemma 2.16 shows that $K_{0}$ is a finite projective $A$-module. At this point, we have
@@ 4.1. Isocrystal from a descent datum
$F_{M,2}\colon\pi_{2}^{\ast}M\xrightarrow{\varphi^{-1}}\pi_{1}^{\ast}M\xrightarrow{\operatorname{id}_{M}\otimes F_{2}}\pi_{1}^{\ast}M\xrightarrow{\varphi}\pi_{2}^{\ast}M.$
@@
$M{\pi_{2}^{\ast}M}$$\rho_{3}^{\ast}M$$M{\pi_{2}^{\ast}M}$$\rho_{3}^{\ast}M.$$\scriptstyle{\pi_{2}^{\ast}}$$\scriptstyle{F_{M}}$$\scriptstyle{\pi_{13}^{\ast}}$$\scriptstyle{\pi_{23}^{\ast}}$$\scriptstyle{F_{M,2}}$$\scriptstyle{F_{M,3}}$$\scriptstyle{\pi_{2}^{\ast}}$$\scriptstyle{\pi_{13}^{\ast}}$$\scriptstyle{\pi_{23}^{\ast}}$
Both rows are equalizers, and all horizontal maps are $A(\!(t^{\mathbb{R}})\!)$-linear while the vertical maps are $F$-semilinear, if we use the ring homomorphisms $\pi_{2}^{\ast}\colon A(\!(t^{\mathbb{R}})\!)\to A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$ and $\rho_{3}^{\ast}\colon A(\!(t^{\mathbb{R}})\!)\to A(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!)$ to give every module an $A(\!(t^{\mathbb{R}})\!)$-module structure. Therefore we obtain an $F$-semilinear bijection $F_{M}\colon M\to M$, and this defines the functor
@@
$M{\pi_{2}^{\ast}M}$$\pi_{2}^{\ast}M$$\pi_{2}^{\ast}M$$\rho_{3}^{\ast}M$$\rho_{3}^{\ast}M$$\scriptstyle{\varphi\circ\pi_{1}^{\ast}}$$\scriptstyle{\varphi\circ\pi_{1}^{\ast}}$$\scriptstyle{\operatorname{id}}$$\scriptstyle{F_{M}\otimes F_{2}}$$\scriptstyle{\pi_{13}^{\ast}}$$\scriptstyle{\pi_{13}^{\ast}}$$\scriptstyle{\pi_{23}^{\ast}\varphi\circ\pi_{12}^{\ast}}$$\scriptstyle{\operatorname{id}}$$\scriptstyle{F_{M}\otimes F_{3}}$
@@
$M{\pi_{1}^{\ast}M}$$\pi_{1}^{\ast}M$$\scriptstyle{\pi_{1}^{\ast}}$$\scriptstyle{\operatorname{id}}$$\scriptstyle{\operatorname{id}_{M}\otimes F_{2}}$
@@
$A(\!(t^{\mathbb{R}})\!)$$A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$$A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$$\scriptstyle{\pi_{1}^{\ast}}$$\scriptstyle{\operatorname{id}}$$\scriptstyle{F_{2}}$
@@
$M{\pi_{2}^{\ast}M}$$N{\pi_{2}^{\ast}N}$$\pi_{2}^{\ast}M$$\rho_{3}^{\ast}M$$\pi_{2}^{\ast}N$$\rho_{3}^{\ast}N.$$\scriptstyle{\varphi_{M}\circ\pi_{1}^{\ast}}$$\scriptstyle{\varphi_{M}\circ\pi_{1}^{\ast}}$$\scriptstyle{g}$$\scriptstyle{\pi_{13}^{\ast}}$$\scriptstyle{\pi_{2}^{\ast}f}$$\scriptstyle{\varphi_{N}\circ\pi_{1}^{\ast}}$$\scriptstyle{\varphi_{N}\circ\pi_{1}^{\ast}}$$\scriptstyle{\pi_{13}^{\ast}}$$\scriptstyle{\pi_{23}^{\ast}\varphi_{M}\circ\pi_{12}^{\ast}}$$\scriptstyle{h}$$\scriptstyle{\rho_{3}^{\ast}f}$$\scriptstyle{\pi_{23}^{\ast}\varphi_{N}\circ\pi_{12}^{\ast}}$
@@ 4.2. The case of a reduced ring
with the property that each associated graded $M_{i}/M_{i-1}$ is isomorphic to $(k(\!(t^{\mathbb{R}})\!),c_{i}F)$ for some $c_{i}\in k^{\times}$.
@@
has an $i$-dimensional kernel. Since the kernel when restricted to $\pi_{2}^{\ast}M_{i-1}$ is (i1)(i-1)-dimensional, there exist $x_{1},\dotsc,x_{i}\in k(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$ with $x_{i}\neq 0$ satisfying
@@
$\pi_{1}^{\ast}M\otimes_{A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$$M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$$\pi_{2}^{\ast}M\otimes_{A(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$$M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!).$$\scriptstyle{(\prod_{i}\pi_{1})^{\ast}f}$$\scriptstyle{\varphi\otimes\operatorname{id}}$$\scriptstyle{(\prod_{i}\pi_{2})^{\ast}f}$
@@
For every element $\alpha=(\alpha_{i})\in M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$, define the support of $\alpha$ as $\operatorname{supp}(\alpha)=\bigcup_{i\in I}\operatorname{supp}(\alpha_{i})\subseteq\mathbb{R}^{2}$, where the support of $\alpha_{i}\in M_{0,i}(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$ is as in Definition 2.11. For $\beta=(\beta_{i})\in M_{0}\otimes_{A}\prod_{i\in I}k_{i}(\!(t^{\mathbb{R}})\!)$ we similarly define $\operatorname{supp}(\beta)=\bigcup_{i}\operatorname{supp}(\beta_{i})\subseteq\mathbb{R}^{2}$.
@@
inside $\mathbb{R}^{2}$, where by ++ we mean the Minkowski sum. By our assumptions, the right hand side is always contained in $\{(x,y):2x+y\geq C+D_{i}\}$. On the other hand, for some $1\leq j\leq n$, the left hand side contains $(D_{i},0)$. This shows that $2D_{i}\geq C+D_{i}$, and hence $D_{i}\geq C$. This shows that $\operatorname{supp}(m_{j,i})\subseteq[C,\infty)$ for all $i,j$, and taking the union over all $i$, we obtain $\operatorname{supp}(m_{j})\subseteq[C,\infty)$ for all $1\leq j\leq n$. For an arbitrary $m\in M$, we can now write $m=\sum_{j=1}^{n}a_{j}m_{j}$ for $a_{j}\in A(\!(t^{\mathbb{R}})\!)$, and since each $\operatorname{supp}(a_{j})$ is bounded below, we conclude that $\operatorname{supp}(m)$ is bounded below as well.
@@ 4.3. The case of a general ring
The second Čech cohomology group of the quasi-coherent sheaf $\operatorname{Hom}_{A/I}(\bar{M},I\otimes_{A/I}\bar{M})$ with respect to $\{U_{i}\}$ vanishes on $\operatorname{Spec}A/I$ as each $U_{i}$ is affine. Therefore the $2$-cocycle is a coboundary, and hence we may modify the isomorphisms $M_{i}|_{U_{i}\cap U_{j}}\cong M_{j}|_{U_{i}\cap U_{j}}$ so that they do satisfy the cocycle condition. ∎
@@
$\operatorname{Hom}_{A(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!)}(M_{0}(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!),(IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!))$ $\cong\operatorname{Hom}_{A}(M_{0},IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!)$
by Lemma 2.8. Using this isomorphism, we identify $\psi_{+}$ as an element of $\operatorname{Hom}_{A}(M_{0},IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}})\!)$, and similarly the additive cocycle condition as an equation valued in the group $\operatorname{Hom}_{A}(M_{0},IM_{0})(\!(t^{\mathbb{R}},u^{\mathbb{R}},v^{\mathbb{R}})\!)$.
We now note that the cocycle condition implies
$\{(x,0,z):(x,z)$ $\in\operatorname{supp}(\psi_{+})\}=\operatorname{supp}(\pi_{13}^{\ast}\psi_{+})$
$\subseteq\operatorname{supp}(\pi_{23}^{\ast}\psi_{+})\cup\operatorname{supp}(\pi_{12}^{\ast}\psi_{+})\subseteq(\{0\}\times\mathbb{R}^{2})\cup(\mathbb{R}^{2}\times\{0\}).$
It follows $\operatorname{supp}(\psi_{+})$ is contained in $(\{0\}\times\mathbb{R})\cup(\mathbb{R}\times\{0\})$. This allows us to write
$\psi_{+}=c+\sum_{0\neq d\in\mathbb{R}}(a_{d}t^{a}+b_{d}u^{d}),\quad c,a_{d},b_{d}\in\operatorname{Hom}_{A}(M_{0},IM_{0}).$
The cocycle condition then can be written out as
$c+\sum_{0\neq d\in\mathbb{R}}(a_{d}t^{d}+b_{d}v^{d})=2c+\sum_{0\neq d\in\mathbb{R}}(a_{d}t^{d}+(b_{d}+a_{d})u^{d}+b_{d}v^{d})$
and hence $c=0$ and $b_{d}=-a_{d}$ for all $0\neq d\in\mathbb{R}$. Setting $\xi_{+}=\sum_{0\neq d\in\mathbb{R}}a_{d}t^{d}$, we obtain
@@
Regarding $\xi_{+}$ as an $A(\!(t^{\mathbb{R}})\!)$-linear map $M_{0}(\!(t^{\mathbb{R}})\!)\to(IM_{0})(\!(t^{\mathbb{R}})\!)$, we now define $\xi=\operatorname{id}+\xi_{+}$. It follows that